OpenStudy (anonymous):

Simplify: Radical 15/10

OpenStudy (allank):

\[\sqrt{3/2}\]

OpenStudy (anonymous):

how? please explain..

OpenStudy (allank):

Whatever is under the square root sign can be simplified as usual. 15/10 = 3/2

OpenStudy (allank):

Are you okay now?